OpenStudy (anonymous):

A number is between 850 and 900. The number is divisible by 9 and the units digit is 1. What is the number? 801 891 851 871

OpenStudy (anonymous):

Add the digits of the numbers and see that the sum is divisible by 9...Then the number is also divisible by 9.

OpenStudy (anonymous):

Well the easiest way is to guess and check.

OpenStudy (anonymous):

@AbhimanyuPudi : I like that!

OpenStudy (anonymous):

@Dido525 Thanks:) Its easier, isn't it?

OpenStudy (anonymous):

Don't bother guessing the first response... it isn't in the right range of 850-900

OpenStudy (anonymous):

I agree!! @AbhimanyuPudi

OpenStudy (anonymous):

im not getting the answer?

OpenStudy (anonymous):

It is one of those nasty distractor answers they throw in first to try to trick you... it is divisible by 9 and has ones digit of 1, but it's not in the correct range.

OpenStudy (anonymous):

You will get 891 only satisfying that condition.. 8+9+1=18 which is divisible by 9...18/9 = 2

OpenStudy (amorfide):

as @AbhimanyuPudi said, look at the possible choices add up each number so for example 801 8+0+1=9 now divide that by 9 if you can divide it by 9, then the number 801 is also divisible by 9 but that is not your answer since it is less than 850

OpenStudy (anonymous):

Well it's obviously not the first since it's not in the range: 8+9+1=18 . is 18 divisible by 9? 8+5+1=15. Is 15 divisible by 9? 8+7+1=16. Is 16 divisible by 9?